OpenStudy (anonymous):

HELP PLEASE: Is the solution to Amplitude (1/2), period 6pi, horizontal shift two units to the right 1/2 sin 3.28t +2 ?

OpenStudy (jdoe0001):

hmm what's the 3.28? and did you mean \(\bf \cfrac{1}{2}sin(3.28t)+2\quad ?\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3.28 is for the 2pi/6pi which is my desired amount

OpenStudy (ybarrap):

\(A\times sin(\cfrac{2\pi}{T}-\phi)\), where \(A\) is amplitude, \(T\) is period and \(\phi \) is phase shift. You have \(A=\cfrac{1}{2}\), \(T=6\pi\) and \(\phi=2\). Note: \(\cfrac{2\pi}{6\pi}=\cfrac{1}{3}\) for your frequency.

OpenStudy (anonymous):

so the equation would be 1/2 sin(1/3) +2 ?

OpenStudy (ybarrap):

Almost, I forgot the \(t\): $$ \cfrac{1}{2}\times sin(\cfrac{t}{3}-2) $$ Don't forget to include the 2 inside the parenthesis, which shifts the entire sine function to the right by 2 time units.

OpenStudy (anonymous):

thank you

OpenStudy (ybarrap):

you're welcome